If I make 10 cents on every barrel of salt water, and there are 33,405 barrels of salt water, how many dollars do I get? $3340.50 or $334.05?
Answer by jim_thompson5910(35256)

(33,405 barrels )*(10 cents/1 barrel) = 33,405*10 = 334,050 cents

(334,050 cents)*(1 dollar/100 cents) = 334,050/100 = 3340.50 dollars

So the answer is $3340.50

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------